# Generating metamers of human scene understanding

### v2

@@ Abstract
Human vision combines low-resolution “gist” information from the visual periphery with sparse but high-resolution information from fixated locations to construct a coherent understanding of a visual scene. In this paper, we introduce MetamerGen, a tool for generating scenes that are aligned with latent human scene representations. MetamerGen is a latent diffusion model that combines peripherally obtained scene gist information with information obtained from scene-viewing fixations to generate image metamers for what humans understand after viewing a scene. Generating images from both high and low resolution (i.e. “foveated”) inputs constitutes a novel image-to-image synthesis problem, which we tackle by introducing a dual-stream representation of the foveated scenes consisting of DINOv2 tokens that fuse detailed features from fixated areas with peripherally degraded features capturing scene context. To evaluate the perceptual alignment of MetamerGen generated images to latent human scene representations, we conducted a same-different behavioral experiment where participants were asked for a “same” or “different” response between the generated and the original image. With that, we identify scene generations that are indeed metamers for the latent scene representations formed by the viewers. MetamerGen is a powerful tool for understanding scene understanding. Our proof-of-concept analyses uncovered specific features at multiple levels of visual processing that contributed to human judgments. While it can generate metamers even conditioned on random fixations, we find that high-level semantic alignment most strongly predicts metamerism when the generated scenes are conditioned on viewers’ own fixated regions.
@@ 1 Introduction
Inspired by these previous studies showing that generated textures and shapes can become metamers for human scene perception, we introduce MetamerGen, a state-of-the-art generative model that extends the metamer generation approach to human scene understanding. Rather than seeking to generate simple patterns that share low-level statistics with peripheral vision, MetamerGen better captures a post-gist level of representation reflecting multiple free-viewing fixations. We see this topic as closer to scene understanding, and we see the scenes generated by MetamerGen being hypotheses for what people believe to be in their peripheral vision. Henceforth, we will therefore use the term scene metamer to refer to two scenes that have an equivalent understanding. Our approach combines a gist-level scene representation extracted from peripherally blurred pixels with higher-resolution and fixation-specific “foveal” representations corresponding to scene-viewing fixations. Scene gist and the objects fixated during viewing are used to generate in the non-fixated blurred pixels a scene context that is aligned with what a human understands to be in their peripheral vision.
We not only show that many of the scenes generated by MetamerGen are metamers for human scene understanding, we also model the dynamic evolution of this understanding by leveraging the capability of a latent diffusion model (Rombach et al., 2022) to generate photorealistic images from diverse conditioning signals (Sohl-Dickstein et al., 2015; Zhang et al., 2023; Ramesh et al., 2022). Because MetamerGen is a latent diffusion model (Stable Diffusion; Rombach et al., 2022), we can use each viewing fixation as a conditioning signal to obtain an incremental fixation-by-fixation understanding of a scene (Figure 1).
@@
Our conditioning mechanism allows plausible scene hypotheses to be generated from variable input information. We expect more foveal glimpses of a scene will lead to a richer DINOv2 representation that will in turn enable MetamerGen to generate increasingly plausible and contextually appropriate content at the non-fixated scene locations, analogous to how human scene understanding becomes more elaborated with more viewing fixations. We see MetamerGen as a tool for generating fixation-specific scene understanding hypotheses that cognitive scientists can test in behavioral studies.
@@ 2 Preliminaries
2.1 Image generation using latent generative models
@@ 3.1 Representing foveal & peripheral visual features
DINOv2 processes $448\times 448$ images with a patch size of $14\times 14$, yielding $1024$ tokens ($32\times 32$ grid), each embedded in $768$ dimensions (along with a CLS token representing and four register tokens encoding general information about the image). The patch token at a specific location encodes detailed visual and semantic information about that location, analogous to the high-resolution information sampled by the fovea during a fixation. It also encodes local contextual information from around the fixated locations, analogous to the low-resolution parafoveal information extracted by humans(Adeli et al., 2023; 2025). To model the information gathered during a series of fixations, we apply a binary mask $M_{\text{fixation}}$ to the patch tokens extracted from a scene image $I$, corresponding to the image locations fixated by humans, zeroing out all non-fixated image patches.
@@ 3.2 Foveal & peripheral conditioning adapters
Both foveal and peripheral DINOv2 embeddings are first processed through separate Perceiver-based resampler networks $R(\cdot)$ (Alayrac et al., 2022; Jaegle et al., 2021) that compress the 1024 DINOv2 embeddings into 32 conditioning tokens compatible with the pre-trained UNet’s cross-attention (see Appendix A.3 for more information).
@@
which we then combine additively into the denoising through cross-attention.
@@ 3.3 Training and Inference
We start from a pre-trained Stable Diffusion 1.5 network (Rombach et al., 2022). The trainable components of MetamerGen are the foveal and peripheral resampler networks and their associated key-value projection matrices. Training is conducted on the complete MS-COCO training set (Lin et al., 2015) of approximately $118,000$ images. For foveal conditioning, we apply binary masks that randomly retain $\{1,2,3,5,10\}$ DINOv2 patch tokens while zeroing all others. This sampling strategy ensures compatibility with our free-viewing behavioral experiments, which constrain scene viewing to a maximum of 10 fixations. For peripheral conditioning, we blur the images by downsampling to $\{0.0625\times,0.125\times,0.25\times,0.5\times,1\times\}$ of the original resolution.
@@
Figure 2: FID values for different input parameters of MetamerGen. Lower FID values indicate closer alignment with real images and better quality.
@@ 3.4 Image generation quality
We first evaluate the image quality of samples from our model using Fréchet Inception Distance (FID; Heusel et al., 2017) between images generated from MetamerGen and COCO-10k-test. Figure 2 shows the results using a single central fixation. Green: We fix the blur level to $0.25\times$, matching our behavioral paradigm, and evaluate how peripheral context affects generation quality by varying the peripheral scale. As peripheral scale increases, FID scores improve as the model better integrates the context coming from the peripheral DINOv2 representations. Red: We evaluated the effect of blur level and found that our model consistently generates plausible scenes for all levels of blur in our evaluation. Blue: We include a text-to-image baseline using SD-1.5 with 10k random captions from the COCO training set. MetamerGen, fine-tuned on the COCO images, consistently outperforms the text-to-image model, proving that we have successfully integrated images of variable resolution into the conditioning mechanism of Stable Diffusion.
@@ 4.1 Probing latent scene representations through same/different judgments
Metamerism, a term initially established in the color vision literature, has since been applied to questions in texture perception and visual crowding to infer underlying perceptual representation (Freeman and Simoncelli, 2011; Balas et al., 2009; Rosenholtz et al., 2012). In the context of scene understanding, metamers offer a unique opportunity to probe what the visual system encodes and retains after viewing complex natural scenes.
Scene perception requires the extraction of meaningful structures from complex inputs. This includes identifying spatial layout, object relations, and global context (Oliva and Torralba, 2006), and is shaped by what the visual system extracts rather than the stimulus itself. When a person views scene A and forms internal representationA, then later sees a different scene B and forms representationB, we can test whether these representations are perceptually aligned. If scene B serves as a metamer to scene A, aligning these internal representations reveals what information the brain has perceived and retained from the original scene. By using metamerism as a proxy for the content of scene representations, we can systematically investigate the structure of human scene understanding.
@@ 4.2 Real-time behavioral paradigm
Participants (n=45) engaged in a two-stage paradigm that consisted of freely viewing a scene followed by making a same-different judgment about that scene. Each trial followed the same structured sequence (Appendix  A.1.2). In Stage 1, participants completed a drift check and then freely viewed a natural scene image until reaching a predetermined fixation count $\{1,2,3,5,10\}$, after which the image was removed. We systematically varied information availability by manipulating fixation count, testing how additional visual information influenced the generation quality and the rate of metamerism. Critically, participants chose their own fixation locations (see Appendix A.1.1). Stage 2 immediately followed scene viewing offset and began with participants maintaining central fixation throughout a 5-second interval. During this interval, MetamerGen generated in real time a new version of the original image based on the viewing fixations. Participants were then shown a second image for 200 milliseconds—too brief for an eye movement but sufficient for a perceptual judgment (Broderick et al., 2023; Wallis et al., 2019), and asked to indicate whether this second image was the same or different from the image viewed originally. Although our primary experimental manipulation was whether the second image was the original or a generation from the participant’s own fixations, we also had a “random” condition consisting of 12 other participants making a similar same or different judgment, only now between a just-viewed original image and a generation based on a random sampling of fixations in the image. We found that, although random generations did sometimes elicit “same” judgments, generations based on a viewer’s own fixations produced the best estimate of their scene understanding.
Figure 3: Metameric vs. non-metameric judgments. (Left) Original images with human fixations overlaid in red paired with generated images judged as the “same” by participants. (Right) Original images with fixations and generated images judged as “different” by participants. More examples, including generations from random fixations, can be found in Appendix A.6
@@
Our stimuli were 300 images from the YFCC100M dataset (Thomee et al., 2016), a subset of the Visual Genome dataset (Krishna et al., 2017). We chose this dataset to avoid overlap with the COCO dataset used in MetamerGen training. To maximize visual diversity in our 300 image sample, we used DreamSim (Fu et al., 2023) to cluster YFCC100M images in semantic representational space and selected one representative image per cluster. Images were filtered to exclude challenging elements for current diffusion models (e.g., humans, printed text, clocks).
@@ 5 Results: Probing the features of scene metamers
As shown in Figure 3, MetamerGen’s conditioning on a viewer’s own fixations enables it to generate rich and plausible hypotheses for a person’s scene understanding. However, from such qualitative results it is difficult to know the features responsible for a generated scene becoming a metamer. We conducted three quantitative analyses to answer this question. First, we compared visual similarity from neurally grounded CNN features with human same–different responses. Our second analysis focused on interpretability by asking which visual features—ranging from low to high levels—shape metameric judgments. Third, we determined the contribution of foveal versus peripheral information by ablating each from the combined model and observing the effect on metamerism rates.
@@ 5.1 Neurally-grounded feature maps
To investigate the level of visual processing in the human brain that is most responsible for a scene becoming a metamer, we compared our behavioral judgments to an AlexNet that was trained to be robust to image blur and demonstrated to have internal representations aligned with neural responses from visual brain areas ranging from V1 to inferotemporal cortex (IT) (Jang and Tong, 2024). Using this neurally-grounded model, we isolated contributions from early to late levels of the visual hierarchy to metameric perception. As illustrated in Figure 4, our analysis pipeline treats early, middle, and late layers as proxies for different stages of visual processing. For each layer, we extracted feature maps from both original and generated images and computed cosine similarity to quantify alignment across the visual hierarchy. For attention-guided generations, we found that as feature similarity increased, the proportion of participants judging images as metameric also increased. This relationship held consistently across all layers of the network, from early visual features through high-level representations. These results suggest that metamerism spans the entire hierarchy of visual processing rather than being confined to a single processing stage, and that in order for a scene to be judged a metamer there must be broad representational alignment across low-level and high-level visual features.
@@ 5.2.1 Low-level visual features
We compared human “same” judgments as a function of (i) Gabor filter intensities and (ii) Sobel edge density response differences between the generated and original images to assess how low-level texture features affect similarity judgments. We used normalized responses from Gabor filters oriente at 0°, 45°, 90°, and 135°. Surprisingly, we found that positive differences in Gabor filter responses—where generated images showed stronger texture responses than originals—correlated with more “same” judgments. This suggests that enhanced texture definition, which makes boundaries more distinctive, increases the perceived realism of generated images, even when they differ substantially from the originals (Ho et al., 2012). We also found that greater Sobel edge density responses (Kanopoulos et al., 1988) led to greater “same” judgments, though this effect was redundant with the Gabor filter effect (see A.4).
@@ 5.2.2 Mid-level visual features
We tested two different mid-level visual features that represent local scene layout information prior to full scene segmentation: (i) relative depth and (ii) proto-object segmentation.
We used Depth Anything (Yang et al., 2024a) to obtain depth maps from the original and generated images and compared them using Scale-Invariant Log error (SiLog) (Lee et al., 2021; Eigen et al., 2014). As shown in the leftmost plot from Figure 5, as discrepancies between depth maps increased, the proportion of “same” metameric judgments systematically decreased. This pronounced effect of depth on metamerism rate suggests that a mid-level representation of the depth relationships in a scene, essential for capturing scene layout (Verhoef et al., 2016), is an important factor in determining whether a generated scene becomes a metamer.
To analyze the effect of mid-level grouping on metamer rates, we extracted proto-object segmentations using the conv3 layer of the blur-trained AlexNet (Jang and Tong, 2024). These mid-layer representations are important to the formation of robust “proto-objects” (Finkel and Sajda, 1992; Yu et al., 2014), which are the pre-semantic building blocks of visual objects created by grouping local features into simple shapes. We found that greater mIoU in proto-object segmentation predicted more “same” judgments (Figure 5), demonstrating that similarity in proto-object segmentation also affects scene metamerism. Of the two mid-level features tested, however, the effect of depth seemed the clearer.

### v3

@@ Abstract
Human vision combines low-resolution “gist” information from the visual periphery with sparse but high-resolution information from fixated locations to construct a coherent understanding of a visual scene. In this paper, we introduce MetamerGen, a tool for generating scenes that are aligned with latent human scene representations. MetamerGen is a latent diffusion model that combines peripherally obtained scene gist information with information obtained from scene-viewing fixations to generate image metamers for what humans understand after viewing a scene. Generating images from both high and low resolution (i.e. “foveated”) inputs constitutes a novel image-to-image synthesis problem, which we tackle by introducing a dual-stream representation of the foveated scenes consisting of DINOv2 tokens that fuse detailed features from fixated areas with peripherally degraded features capturing scene context. To evaluate the perceptual alignment of MetamerGen generated images to latent human scene representations, we conducted a same-different behavioral experiment where participants were asked for a “same” or “different” response between the generated and the original image. With that, we identify scene generations that are indeed metamers for the latent scene representations formed by the viewers. MetamerGen is a powerful tool for understanding scene understanding. Our proof-of-concept analyses uncovered specific features at multiple levels of visual processing that contributed to human judgments. While it can generate metamers even conditioned on random fixations, we find that high-level semantic alignment most strongly predicts metamerism when the generated scenes are conditioned on viewers’ own fixated regions. [^1] Project Link: https://rainarit.github.io/metamergen/
@@ 1 Introduction
Inspired by these previous studies showing that generated textures and shapes can become metamers for human scene perception, we introduce MetamerGen, a state-of-the-art generative model that extends the metamer generation approach to human scene understanding. Rather than seeking to generate simple patterns that share low-level statistics with peripheral vision, MetamerGen better captures a post-gist level of representation reflecting multiple free-viewing fixations. We see this topic as closer to scene understanding, and we consider the scenes generated by MetamerGen as hypotheses for what people believe to be in their peripheral vision. In this paper, we will therefore use the term scene metamer to refer to two scenes that have an equivalent understanding. Our approach combines a gist-level scene representation extracted from peripherally blurred pixels with higher-resolution and fixation-specific “foveal” representations corresponding to scene-viewing fixations. Scene gist and the objects fixated during viewing are used to generate in the non-fixated blurred pixels a scene context that is aligned with what a human understands to be in their peripheral vision.
We not only show that many of the scenes generated by MetamerGen are metamers for human scene understanding, we also model the dynamic evolution of this understanding by leveraging the capability of a latent diffusion model (Rombach et al., 2022) to generate photorealistic images from diverse conditioning signals (Zhang et al., 2023; Ramesh et al., 2022). MetamerGen is a latent diffusion model (Stable Diffusion; Rombach et al., 2022), thus we use each viewing fixation as a conditioning signal to obtain an incremental fixation-by-fixation understanding of a scene (Figure 1).
@@
Our conditioning mechanism allows plausible scene hypotheses to be generated from variable input information. We expect more foveal glimpses of a scene will lead to a richer DINOv2 representation that will in turn enable MetamerGen to generate increasingly plausible and contextually appropriate content at the non-fixated scene locations, analogous to how human scene understanding becomes more elaborate with more viewing fixations. We see MetamerGen as a tool for generating fixation-specific scene understanding hypotheses that cognitive scientists can test in behavioral studies.
@@ 2 Preliminaries
2.1 Image generation using latent diffusion models
@@ 3.1 Representing foveal & peripheral visual features
DINOv2 processes $448\times 448$ images with a patch size of $14\times 14$, yielding $1024$ tokens ($32\times 32$ grid), each embedded in $768$ dimensions (along with a CLS token and four register tokens encoding general information about the image). The patch token at a specific location encodes detailed visual and semantic information about that location, analogous to the high-resolution information sampled by the fovea during a fixation. It also encodes local contextual information from around the fixated locations (Adeli et al., 2023; 2025), analogous to the low-resolution parafoveal information extracted by humans. To model the information gathered during a series of fixations, we apply a binary mask $M_{\text{fixation}}$ to the patch tokens extracted from a scene image $I$, corresponding to the image locations fixated by humans, zeroing out all non-fixated image patches.
@@ 3.2 Foveal & peripheral conditioning adapters
Both foveal and peripheral DINOv2 embeddings are first processed through separate Perceiver-based resampler networks $R(\cdot)$ (Alayrac et al., 2022; Jaegle et al., 2021) that compress the 1024 DINOv2 embeddings into 32 conditioning tokens compatible with the pre-trained UNet’s cross-attention (see Appendix A.3 for more information):
@@
which we then combine additively into the denoising through cross-attention:
@@ 3.3 Training and Inference
We start from a pre-trained Stable Diffusion 1.5 network (Rombach et al., 2022). The trainable components of MetamerGen are the foveal and peripheral resampler networks and their associated key-value projection matrices. Training is conducted on the complete MS-COCO training set (Lin et al., 2015) of approximately $118,000$ images. For foveal conditioning, we apply binary masks that randomly retain $\{1,2,3,5,10\}$ DINOv2 patch tokens while zeroing all others. This sampling strategy ensures compatibility with our free-viewing behavioral experiments, which constrain scene viewing to a maximum of 10 fixations. At training time, the foveal locations are randomly sampled, whereas at inference time, they are determined by the experimental design. For peripheral conditioning, we blur the images by downsampling to $\{0.0625\times,0.125\times,0.25\times,0.5\times,1\times\}$ of the original resolution.
@@
Figure 2: FID values for different input parameters of MetamerGen. Lower FID indicates closer alignment with real images and better quality.
@@ 3.4 Image generation quality
We first evaluate the image quality of samples from our model using Fréchet Inception Distance (FID; Heusel et al., 2017) between images generated from MetamerGen and COCO-10k-test. Figure 2 shows the results using a single central fixation. Green: We fix the blur level to $0.25\times$, matching the blur level used in our behavioral paradigm, and evaluate how peripheral context affects generation quality by varying the peripheral scale. As peripheral scale increases, FID scores improve as the model better integrates the context coming from the peripheral DINOv2 representations. Red: We evaluated the effect of blur level and found that our model consistently generates plausible scenes for all levels of blur in our evaluation. Blue: We include a text-to-image baseline using SD-1.5 with 10k random captions from the COCO training set. MetamerGen, fine-tuned on the COCO images, consistently outperforms the text-to-image model, proving that we have successfully integrated images of variable resolution into the conditioning mechanism of Stable Diffusion.
@@ 4.1 Probing latent scene representations through same/different judgments
Metamerism, a term initially established in the color vision literature, has since been applied to questions in texture perception and visual crowding to infer underlying perceptual representations (Freeman and Simoncelli, 2011; Balas et al., 2009; Rosenholtz et al., 2012). In the context of scene understanding, metamers offer a unique opportunity to probe what the visual system encodes and retains after viewing complex natural scenes.
Scene understanding requires the extraction of meaningful structures from complex inputs, including spatial layout, object relations, and global context (Oliva and Torralba, 2006). It is shaped by what the visual system extracts from the input and not the input itself. When a person views scene A and forms internal representationA, then later sees a different scene B and forms representationB, they may believe that these scenes are the same if these representations are aligned. We infer aligned representations when a person responds “same” to a generated version of the scene. These scene metamers therefore reveal what information the brain has encoded and retained from an originally viewed scene, and by exploring the factors causing a generation to become a metamer we are able to investigate the structure of human scene understanding.
@@ 4.2 Real-time behavioral paradigm
Participants (n=45) engaged in a paradigm requiring them to first freely view a scene and then make a same-different judgment about that scene. Each of the 300 experimental trials followed the same structured sequence, and eye gaze was tracked throughout each trial (see Appendix  A.1 for additional details and a figure illustrating the paradigm). Participants completed a drift check (to maintain eye-data quality) and then freely viewed a natural scene image until reaching a predetermined fixation count $\{1,2,3,5,10\}$, after which the image was removed. Critically, participants chose their own fixation locations. We systematically varied information availability by manipulating fixation count, testing how additional visual information influenced the generation quality and the rate of metamerism. A 5-second interval followed the offset of the scene during which the participant maintained fixation on an otherwise blank screen. During this interval, MetamerGen generated a new version of the original image in real time based on the viewing fixations. Participants were then shown a second image for 200 milliseconds, too brief for an eye movement but sufficient for a perceptual judgment (Broderick et al., 2023; Wallis et al., 2019), and asked to indicate whether this second image was the same or different from the image viewed originally. Although our primary experimental manipulation was whether the second image was the original or a generation from the participant’s own fixations, we also had a “random” condition consisting of 12 other participants making a similar same or different judgment, only now between a just-viewed original image and a generation based on a random sampling of fixations in the image.
Figure 3: Metameric vs. non-metameric judgments. (Left) Original images with human fixations overlaid in red paired with generated images judged as the “same” by participants. (Right) Original images with fixations and generated images judged as “different” by participants. More examples, including generations from random fixations, can be found in Appendix A.6.
@@
Our stimuli were 300 images from the Visual Genome dataset (Krishna et al., 2017), a subset of the YFCC100M dataset (Thomee et al., 2016). We chose this dataset to avoid overlap with the COCO data used in MetamerGen training. To maximize visual diversity in our 300 image sample, we used DreamSim (Fu et al., 2023) to cluster Visual Genome images in semantic representational space and selected one representative image per cluster. Images were filtered to exclude challenging elements for current diffusion models (e.g., humans, printed text, clocks).
@@ 5 Results: Probing the features of scene metamers
As shown in Figure 3, MetamerGen’s conditioning on a viewer’s own fixations enables it to generate rich and plausible hypotheses for a person’s scene understanding. However, from such qualitative results it is difficult to know the features responsible for a generated scene becoming a metamer. We conducted three quantitative analyses to answer this question. First, we compared visual similarity from neurally grounded CNN features with human same–different responses. Our second analysis focused on interpretability by asking which visual features—ranging from low to high levels—shape metameric judgments. Third, we collected new behavioral data to determine the contribution of foveal versus peripheral information by ablating each from the combined model and observing the effect on metamerism rates.
@@ 5.1 Neurally-grounded feature maps
To investigate the level of visual processing in the human brain that is most responsible for a scene becoming a metamer, we compared our behavioral judgments to an AlexNet that was trained to be robust to image blur and demonstrated to have internal representations aligned with neural responses from visual brain areas ranging from V1 to inferotemporal cortex (IT) (Jang and Tong, 2024). Using this neurally-grounded model, we isolated contributions from early to late levels of the visual hierarchy to metameric perception. As illustrated in Figure 4, our analysis pipeline treats early, middle, and late layers as proxies for different stages of visual processing. For each layer, we extracted feature maps from both original and generated images and computed cosine similarity to quantify alignment across the visual hierarchy. For attention-guided generations, we found that as feature similarity increased, the proportion of participants judging images as metameric also increased. This relationship held consistently across all layers of the network, from early visual features through high-level representations. These results suggest that metamerism spans the entire hierarchy of visual processing rather than being confined to a single processing stage, and that in order for a scene to be judged a metamer, there must be broad representational alignment across low and high-level visual features.
@@ 5.2.1 Low-level visual features
We compared human “same” judgments as a function of (i) Gabor filter intensities and (ii) Sobel edge density response differences between the generated and original images to assess how low-level texture features affect similarity judgments. We used normalized responses from Gabor filters oriented at 0°, 45°, 90°, and 135°. Surprisingly, we found that positive differences in Gabor filter responses—where generated images showed stronger texture responses than originals—correlated with more “same” judgments. This suggests that enhanced texture definition, which makes boundaries more distinctive, increases the perceived realism of generated images, even when they differ substantially from the originals (Ho et al., 2012). We also found that greater Sobel edge density responses (Kanopoulos et al., 1988) led to greater “same” judgments, though this effect was redundant with the Gabor filter effect (see A.5).
@@ 5.2.2 Mid-level visual features
We tested two different mid-level visual features that represent local scene layout information prior to full scene segmentation: (i) relative depth and (ii) proto-object segmentation. We used Depth Anything (Yang et al., 2024a) to obtain depth maps from the original and generated images and compared them using Scale-Invariant Log error (SiLog) (Lee et al., 2021; Eigen et al., 2014). As shown in the leftmost plot from Figure 5, as discrepancies between depth maps increased, the proportion of “same” metameric judgments systematically decreased. This pronounced effect of depth on metamerism rate suggests that a mid-level representation of the depth relationships in a scene, essential for capturing scene layout (Verhoef et al., 2016), is an important factor in determining whether a generated scene becomes a metamer.
To analyze the effect of mid-level grouping on metamer rates, we extracted proto-object segmentations using the conv3 layer of the blur-trained AlexNet (Jang and Tong, 2024). These mid-layer representations are important to the formation of robust “proto-objects” (Finkel and Sajda, 1992; Rensink, 2000; Yu et al., 2014), which are the pre-semantic building blocks of visual objects created by grouping local features into simple shapes. We found that greater mIoU in proto-object segmentation predicted more “same” judgments (Figure 5), demonstrating that similarity in proto-object segmentation also affects scene metamerism. Of the two mid-level features tested, however, the effect of depth seemed the clearer.